OpenStudy (anonymous):

draw the graph of a function 2x-3y=9

OpenStudy (anonymous):

put into slope intercept form. which is y = 2/3x - 3. one point is the y intercept at (0, -3). from that point go up two points, to the right 3 points (from the slope 2/3x) and put a point there. then draw a srtaight line betwene the points.

OpenStudy (agreene):

2x-3y=9 -3y=2x-9 y=2/3x-3 <=== graph that

OpenStudy (anonymous):

We need to solve for y. 2x - 3y = 9 2x - 3y - 2x = 9 - 2x -3y = 9 - 2x Multiply by -1 across -(-3y) = -9 + 2x 3y = -9 + 2x Divide through by 3 y = -3 + 2/3x Rearrange: y = 2/3x - 3 Now plot in values for x. Typically use like x = -5 up to x = 5. This will give you the graph. Hope this helps!

OpenStudy (anonymous):

should i put values for x and also for y?

OpenStudy (anonymous):

Just put in a value for x and that will give you a value for y. Then you have the ordered pair (x,y). For instance take x = 2 y = 2/3(2) - 3 y = 4/3 - 3 y = 4/3 - 9/3 y = -5/3 This gives us: (2,-5/3) We can rewrite y = 2/3x-3 as f(x)=2/3x - 3 This means y is a function of x. Hope this clarifies a bit.